OpenStudy (anonymous):

I will become your fan James added 3 and two-thirds plus 4 and one-eighth. His work is shown here. What did James do right? What did he do wrong? What should he have done? three and two-thirds plus 4 and one-eighth equals 7 and three-elevenths

jigglypuff314 (jigglypuff314):

he added the denominators directly which is not allowed... \[\frac{ 2 }{ 3 } + \frac{ 1 }{ 8 } \neq \frac{ 2+1 }{ 3+8 }\]

OpenStudy (anonymous):

Can you explain it to me more

jigglypuff314 (jigglypuff314):

for example a half of a pizza plug a fourth of a pizza would not equal a sixth of a pizza \[\frac{ 1 }{ 2 } + \frac{ 1 }{ 4 } \neq \frac{ 1 + 1 }{ 2 + 4 } = \frac{ 2 }{ 6 }\] instead it would be \[\frac{ 1 }{ 2 } + \frac{ 1 }{ 4 } = \frac{ 2 }{ 4 } + \frac{ 1 }{ 4 } = \frac{ 2+1 }{ 4 } = \frac{ 3}{ 4 }\]

OpenStudy (anonymous):

I see thank you

jigglypuff314 (jigglypuff314):

remember that fractions can only be added to fractions that have the same denominator :)